A student is asked to determine the value of Ka for nitrous acid by titration with potassium hydroxide.

The student begins titrating a 48.3 mL sample of a 0.360 M aqueous solution of nitrous acid with a 0.216 M aqueous potassium hydroxide solution, but runs out of standardized base before reaching the equivalence point. The student's last observation is that when 53.2 milliliters of potassium hydroxide have been added, the pH is 3.603.

What is Ka for nitrous acid based on the student's data?
Answer #1

The millimoles of nitrous acid = 48.3 x 0.360 = 17.388

The millimoles of KOH added = 53.2 x 0.216 = 11.491

17.388 - 11.491 = 5.897 millimoles acid left

11.491 millimoles salt formed

Total volume = 48.3 + 53.2 = 101.5 mL

[acid] = 5.897 / 101.5 = 0.0580

[salt] = 11.491 / 101.5 = 0.1132

Now,

pH = pKa + log [salt] / [acid]

3.603 = pKa + log [0.1132 / 0.0580]

3.603 = pKa + log (1.951)

3.603 = pKa + 0.2902

pKa = 3.603 - 0.2902

pKa = 3.3128

Ka = 10-3.3128

Ka = 0.000486

= 4.86 * 10-4